# ARCAS: An Augmented Reality Collision Avoidance System with SLAM-Based Tracking for Enhancing VRU Safety

### v1

@@ III-D Automatic Headset Calibration
The system performs automatic calibration by aligning coordinates generated by the 3D LiDAR with those tracked by the headset internal system. This process relies on synchronized 3D position data from both the AR headset and the LiDAR, captured as the user moves within the LiDAR detection range. Using this paired data, the calibration task is formulated as a 3D optimization problem to estimate the rotation matrix $\mathbf{R}$ and translation offset $(X_{0},Y_{0},Z_{0})$, as defined in Equation 9.
$\min_{\mathbf{R},\,X_{0},\,Y_{0},\,Z_{0}}\sum_{i=1}^{n}\left\|\begin{bmatrix}X(i)\\ Y(i)\\ Z(i)\end{bmatrix}-\mathbf{R}\begin{bmatrix}x(i)\\ y(i)\\ z(i)\end{bmatrix}-\begin{bmatrix}X_{0}\\ Y_{0}\\ Z_{0}\end{bmatrix}\right\|^{2}$ (9)
@@
$\min_{\Omega\;\;\text{s.t.}\;\;\Omega^{T}\Omega=I}\left\|\Omega A-B\right\|_{F}$ (10)
where $A$ and $B$ are the centered coordinate matrices, $\Omega$ is the sought 3D rotation matrix, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
In practice, an initial estimate of the linear transformation between the two coordinate sets is obtained, after which a singular value decomposition (SVD) is applied to enforce orthogonality. If the unconstrained transformation is written as $T=U\Sigma V^{T}$, the closest proper 3D rotation matrix in the Frobenius norm sense is given by:
$\mathbf{R}=UV^{T}.$
@@ III-E Off Field-of-View Target Pointer
In real-world scenarios, vehicles or VRUs posing a collision risk may fall outside the headset user field of view (FoV). To address this limitation, the ARCAS incorporates a directional pointer cue to alert the user to off-screen targets approaching or in potential collision path. These cues help the AR headset users identify threats and respond accordingly, even when visual occlusions or limited FoV prevent direct line-of-sight.
To determine whether a target lies outside the FoV, the system computes the angle between the headset user forward-facing direction vector and the vector pointing from the ego user to the moving target. If this angle exceeds a predefined FoV threshold, a directional pointer is activated and displayed at the screen edge to indicate the target’s direction. Otherwise, no pointer is rendered, assuming the user can see the target directly. This approach is inspired by video games, where off-screen objects are represented as markers pointing toward their relative direction in a top-down or first-person view.
@@ III-F Algorithms
This subsection outlines the core logic of the ARCAS system: the server and the client. The server, running on a machine connected to the 3D LiDAR, handles object tracking, and data transmission. The client, running on the AR headset, manages system initialization, coordinate calibration, and real-time visualization. The pseudocode for each component is presented below.
Algorithm 1 Server Pseudocode Logic
@@
Algorithm 2 Client Pseudocode Logic (AR Headset)
@@
Algorithm 3 AR Directional Arrow Display Logic
0: Target distance $d$, relative angle $\theta$ (deg), field of view $\mathrm{FoV}=90^{\circ}$
@@ III-G Multi-Headset Operation
In addition to single-user operation, the ARCAS framework supports multiple AR headsets sharing the same world coordinate frame. This capability is useful for collaborative scenarios such as indoor firefighting or emergency response, where several users must maintain a consistent view of nearby hazards and each other’s positions.
Multi-headset calibration is performed through a simple reference-point procedure. At the start of the session, all users briefly stand at the same physical location and face approximately the same direction. Each headset records its SLAM pose at this point, denoted $\mathbf{p}^{(j)}_{\text{ref}}$ for headset $j$, while one headset is chosen as the primary frame, and the remaining headsets are aligned to it.
For each additional headset $j$, a short calibration motion is performed while both headsets record synchronized 3D poses. Using the paired pose samples $\{\mathbf{p}^{(1)}_{i}\}$ and $\{\mathbf{p}^{(j)}_{i}\}$, the system estimates a 3D rigid-body transformation $(\mathbf{R}_{j1},\mathbf{t}_{j1})$ that maps coordinates from headset $j$ into the primary headset frame, following the same Orthogonal Procrustes formulation used for LiDAR–headset calibration:
$\min_{\mathbf{R}_{j1},\,\mathbf{t}_{j1}}\sum_{i=1}^{n}\left\|\mathbf{p}^{(1)}_{i}-\left(\mathbf{R}_{j1}\,\mathbf{p}^{(j)}_{i}+\mathbf{t}_{j1}\right)\right\|^{2}.$
@@ IV-A Hardware Configuration and Physical Setup
To evaluate the ARCAS system in a real-world setting, several experiments were conducted indoors and outdoors of the Engineering Education and Research Center (EER) at the University of Texas at Austin.
The sensing module consists of a Velodyne VLP-16, a compact 360-degree 3D LiDAR installed at the intersection. The sensor captures motion data in the form of 3D point clouds within a range of up to 100 meters and operates at rotation rates up to 20 Hz, providing a horizontal angular resolution as fine as 0.1° and a vertical resolution of 2.0°. These specifications enable reliable detection of small positional changes with centimeter-level accuracy, supporting real-time tracking of moving targets and continuous updates to their bounding boxes.
The LiDAR was connected to a Dell G15 laptop using a Gigabit Ethernet interface, as the Velodyne VLP-16 transmits 3D point cloud data over UDP via Ethernet connection. The laptop, functioning as the system server, was equipped with an Intel 13th Gen Core i7 processor, 32 GB of RAM, and an NVIDIA GeForce RTX 4060 GPU. This computational setup supported real-time object detection, tracking, and wireless transmission of motion data to the AR headset over a local TCP network.
The AR interface was deployed on a MQPro headset operating in passthrough mode. The headset uses an internal SLAM-based tracking system to determine the user spatial position and orientation. Upon receiving real-time position data from the server, the headset performs coordinate transformations from the world frame to its local reference frame and renders visual cues, such as 3D bounding boxes and directional arrows, directly in the user field of view.
@@
Figure 2: Overview of the pedestrian–e-scooter collision scenario. (a) The top panel presents a side-view illustration alongside the corresponding real-world scene, in which an AR-equipped pedestrian approaches an intersection while an e-scooter rider—without AR and detected solely by 3D LiDAR—emerges from behind an occluding wall. (b) The bottom panel shows the reconstructed trajectories of the pedestrian and the e-scooter rider, including the points at which each reaches its peak velocity.
IV-B Experiments Overview
@@
To assess the effectiveness of ARCAS in enhancing early hazard perception, two real-world outdoor experiments were conducted involving: (1) a pedestrian and an e-scooter rider, and (2) a pedestrian and a vehicle. All experiments took place in a controlled sidewalk and parking-lot area outside the EER building. Each scenario was evaluated under three sensing and visualization conditions: (A) Baseline (Unaided Vision), (B) AR + LiDAR Detection, and (C) AR + Multi-Headset Sharing. For each of the two scenarios, 30 trials were performed under each of the three conditions, resulting in a total of 180 outdoor trials. Time-to-Collision (TTC) was used as the primary metric for evaluating early hazard perception.
@@ IV-C Experimental Scenarios
Scenario 1: Pedestrian–E-scooter Interaction. A pedestrian and an e-scooter rider approach a shared intersection from orthogonal directions. The e-scooter emerges from a fully occluded 90 corner at 10–15 km/h, while the pedestrian walks toward the expected intersection point (EIP) at normal pace as shown in Figure 2.
Scenario 2: Pedestrian–Vehicle Interaction. This scenario involves a pedestrian crossing near a side road while a vehicle approaches from a different direction. The vehicle emerges from an approximately 80 curved corner at 20–30 km/h, gradually entering the pedestrian’s path of motion. Figure 3 presents the layout of this real-world interaction.
@@ IV-D Experimental Conditions
Condition A: Unaided Eye (Baseline). Neither the pedestrian nor the counterpart (e-scooter rider or vehicle driver) wears an AR device. Situational awareness depends entirely on natural human vision. Ground-truth positions and velocities for TTC computation were obtained from VLP-16 LiDAR trajectories.
Condition B: AR + LiDAR Detection. The pedestrian wears a MQPro AR headset, while the counterpart does not. The counterpart’s position and velocity are estimated using the VLP-16 LiDAR, and the pedestrian receives real-time AR cues (3D bounding boxes and directional arrows) corresponding to the detected hazards.
Condition C: AR + Multi-Headset Sharing. Both the pedestrian and the counterpart wear MQPro headsets. SLAM-based multi-headset alignment provides a unified world coordinate frame, enabling direct AR-to-AR exchange of headset poses without relying on LiDAR. Although occasional SLAM resets occurred during fast vehicle motion, the overall tracking remained sufficiently stable for TTC analysis.
IV-E Time-To-Collision (TTC) Computation
@@
$\mathrm{TTC}=\frac{d_{\mathrm{EIP}}}{v}$ (11)
@@
Figure 3: Overview of the pedestrian–vehicle collision scenario. (a) The top panel shows a bird’s-eye illustration alongside the corresponding real-world scene, in which an AR-equipped pedestrian approaches an intersection while a vehicle emerges from an approximately 80 curved and visually occluded side road. (b) The bottom panel illustrates the reconstructed trajectories of both the pedestrian and the vehicle obtained from AR headsets, including the positions at which each reaches its peak velocity.
TABLE I: Time to Collision (TTC) across different sensing and visualization configurations for two real-world near-collision scenarios.
@@ IV-F Summary of Results
Table I presents the average Time-To-Collision (TTC) values computed across 30 repeated trials for each configuration, quantitatively demonstrating that ARCAS substantially improves early hazard perception across both scenarios. Several key findings emerge. Across all conditions, TTC values for the pedestrian increased substantially under AR guidance. In the pedestrian–e-scooter scenario, TTC increased from 2.75 s in the Unaided Eye baseline to 5.51 s with AR + LiDAR and 5.43 s with Multi-Headset Sharing—indicating an approximately twofold improvement in the pedestrian’s available reaction time. The e-scooter rider also exhibited substantial gains, with TTC increasing from 0.69 s to approximately 2.86 s under AR conditions—more than a fourfold improvement. In the pedestrian–vehicle scenario, AR + LiDAR Detection produced the highest pedestrian TTC (6.42 s), while Multi-Headset Sharing delivered a similar improvement (6.29 s) despite occasional SLAM resets. Overall, these results confirm that ARCAS—whether powered by external LiDAR sensing or AR-to-AR SLAM sharing—provides significantly earlier hazard awareness and larger safety margins for VRUs in occluded and fast-approaching traffic conditions.
@@ V Conclusion
This paper introduced ARCAS, a real-time Augmented Reality Collision Avoidance System designed to enhance vulnerable road users’ situational awareness in dynamic traffic environments. By integrating 360-degree 3D LiDAR sensing, SLAM-based headset tracking, and automatic 3D calibration, the system delivers world-locked visual cues—such as 3D bounding boxes and directional arrows—directly in the user’s passthrough view. ARCAS further supports multi-headset alignment, enabling multiple VRUs to share a unified world frame for coordinated awareness.
The system was evaluated through two real-world near-collision scenarios involving pedestrian interactions with an e-scooter rider and a vehicle. Across 180 trials, AR visualization consistently increased Time-To-Collision (TTC) for both the pedestrian and the counterpart. In the pedestrian–e-scooter scenario, the pedestrian’s TTC nearly doubled and the rider’s TTC increased more than fourfold under AR guidance. In the pedestrian–vehicle scenario, LiDAR-based detection yielded the greatest improvements, while multi-headset sharing achieved similar gains despite occasional SLAM resets. These results demonstrate that AR-based visual cues—whether from external sensing or AR-to-AR pose sharing—provide substantially earlier hazard awareness and larger reaction windows in occluded or fast-approaching conditions.
Overall, ARCAS establishes a practical foundation for AR-based safety systems that deliver real-time, personalized hazard awareness for VRUs. Future work will explore intent prediction, adaptive cue generation, and context-aware filtering, as well as large-scale multi-user deployments. Moreover, the designed experiments were conducted in a controlled campus environment, and future work should evaluate ARCAS in more complex and dense urban traffic settings. This future work is supposed to advance intelligent, human-centered safety solutions for next-generation mobility systems.

### v2

@@ III-D Automatic Headset Calibration
The system performs automatic calibration by aligning coordinates generated by the 3D LiDAR with those tracked by the headset internal system. This process relies on synchronized 3D position data from both the AR headset and the LiDAR, captured as the user moves within the LiDAR detection range. Using this paired data, the calibration task is formulated as a 3D optimization problem to estimate the rotation matrix $\mathbf{R}$ and translation offset $(X_{0},Y_{0},Z_{0})$, as defined in Equation (9).
$\min_{\mathbf{R},\,X_{0},\,Y_{0},\,Z_{0}}\sum_{i=1}^{n}\left\|\begin{bmatrix}X(i)\\ Y(i)\\ Z(i)\end{bmatrix}-\mathbf{R}\begin{bmatrix}x(i)\\ y(i)\\ z(i)\end{bmatrix}-\begin{bmatrix}X_{0}\\ Y_{0}\\ Z_{0}\end{bmatrix}\right\|^{2}.$ (9)
@@
$\min_{\Omega\;\;\text{s.t.}\;\;\Omega^{T}\Omega=I}\left\|\Omega A-B\right\|_{F},$ (10)
where $A$ and $B$ are centered coordinate matrices, $\Omega$ is the sought 3D rotation matrix, and $\|\cdot\|_{F}$ is the Frobenius norm.
In practice, an initial estimate of the linear transformation between the two coordinate sets is obtained, after which a singular value decomposition is applied to enforce orthogonality. If the unconstrained transformation is written as $T=U\Sigma V^{T}$, the closest proper 3D rotation matrix in the Frobenius norm sense is given by:
$\mathbf{R}=UV^{T}.$ (11)
@@ III-E Off Field-Of-View Target Pointer
In real-world scenarios, vehicles or VRUs posing a collision risk may fall outside the headset user field of view (FoV). To address this limitation, ARCAS incorporates a directional pointer cue to alert the user to off-screen targets approaching or in potential collision path. These cues help the AR headset users identify threats and respond accordingly, even when visual occlusions or limited FoV prevent direct line of sight. To determine whether a target lies outside the FoV (106°), the system computes the angle between the headset user forward-facing direction vector and the vector pointing from the ego user to the moving target. If this angle exceeds a predefined FoV threshold, a directional pointer is activated and displayed at the screen edge to indicate the target’s direction. Otherwise, no pointer is rendered, assuming the user can see the target directly. This approach is inspired by video games, where off-screen objects are represented as markers pointing toward their relative direction in a top-down or first-person view.
@@ III-F Algorithms
The server, running on a machine connected to the 3D LiDAR, handles object tracking, and data transmission. The client, running on the AR headset, manages system initialization, coordinate calibration, and real-time visualization. The pseudocode for each component is presented below.
@@
Algorithm 1 Server Pseudocode Logic
@@
Algorithm 2 Client Pseudocode Logic (AR Headset)
0: Target distance $d$, relative angle $\theta$ (deg), field of view $\mathrm{FoV}=106^{\circ}$
@@ III-G Multi-Headset Operation
In addition to single user operation, the ARCAS framework supports multiple AR headsets sharing the same world coordinate frame. This capability is useful for collaborative scenarios such as indoor firefighting or emergency response, where several users must maintain a consistent view of nearby hazards and each other’s positions.
Multi-headset calibration is performed through a simple reference point procedure. At the start of the session, all users briefly stand at the same physical location and face approximately the same direction. Each headset records its SLAM pose at this point, denoted $\mathbf{p}^{(j)}_{\text{ref}}$ for headset $j$, while one headset is chosen as the primary frame, and the remaining headsets are aligned to it.
For each additional headset $j$, a short calibration motion is performed while both headsets record synchronized 3D poses. Using the paired pose samples $\{\mathbf{p}^{(1)}_{i}\}$ and $\{\mathbf{p}^{(j)}_{i}\}$, the system estimates a 3D rigid body transformation $(\mathbf{R}_{j1},\mathbf{t}_{j1})$ that maps coordinates from headset $j$ into the primary headset frame, following the same Orthogonal Procrustes formulation used for LiDAR–headset calibration:
$\min_{\mathbf{R}_{j1},\,\mathbf{t}_{j1}}\sum_{i=1}^{n}\left\|\mathbf{p}^{(1)}_{i}-\left(\mathbf{R}_{j1}\,\mathbf{p}^{(j)}_{i}+\mathbf{t}_{j1}\right)\right\|^{2}.$ (12)
@@ IV-A Hardware Configuration and Physical Setup
To evaluate the ARCAS system in a real-world setting, several experiments were conducted indoors and outdoors of the Engineering Education and Research (EER) Center at the University of Texas at Austin. The sensing module consists of a Velodyne VLP-16, a compact 360° 3D LiDAR installed at the intersection. The sensor captures motion data in the form of 3D point clouds within a range of up to 100 meters and operates at rotation rates up to 20 Hz, providing a horizontal angular resolution as fine as 0.1° and a vertical resolution of 2.0°. These specifications enable reliable detection of small positional changes with centimeter level accuracy, supporting real-time tracking of moving targets and continuous updates to their bounding boxes. The LiDAR was connected to a Dell G15 laptop using a Gigabit Ethernet interface, as the Velodyne VLP-16 transmits 3D point cloud data over UDP via Ethernet connection. The laptop, functioning as the system server, was equipped with an Intel 13th Gen Core i7 processor, 32 GB of RAM, and an NVIDIA GeForce RTX 4060 GPU. All point cloud processing, object detection, and Kalman Filter are performed on the CPU, and the dedicated GPU is not required for real-time operation. Based on observed computational loads, minimum requirements for real-time performance at 20 Hz are an Intel Core i7 (9th generation or equivalent) with 8 GB RAM. This setup supported real-time object detection, tracking, and wireless transmission of motion data to the AR headset over a local TCP network. The AR interface was deployed on a MQPro headset operating in passthrough mode. The headset uses an internal SLAM-based tracking system to determine the user spatial position and orientation. Upon receiving real-time position data from the server, the headset performs coordinate transformations from the world frame to its local reference frame and renders visual cues, such as 3D bounding boxes and directional arrows, in the user’s FoV.
@@
Figure 2: Overview of the pedestrian–e-scooter collision scenario. (a) The top panel presents a side view illustration alongside the corresponding real-world scene, in which an AR-equipped pedestrian approaches an intersection while an e-scooter rider, without AR and detected solely by 3D LiDAR, emerges from behind an occluding wall. (b) The bottom panel shows the reconstructed trajectories of the pedestrian and the e-scooter rider, including the points at which each reaches its peak velocity.
Figure 3: Overview of the pedestrian–vehicle collision scenario. (a) The top panel shows a bird’s eye illustration alongside the corresponding real-world scene, in which an AR-equipped pedestrian approaches an intersection while a vehicle emerges from an approximately 80 curved and visually occluded side road. (b) The bottom panel illustrates the reconstructed trajectories of both the pedestrian and the vehicle obtained from AR headsets, including the positions at which each reaches its peak velocity.
IV-B Experiment Overview
@@
To assess the effectiveness of ARCAS in enhancing early hazard perception, two real-world outdoor experiments were conducted involving: (1) a pedestrian and an e-scooter rider, and (2) a pedestrian and a vehicle. All experiments took place in a controlled sidewalk and parking lot area outside the EER building. Each scenario was evaluated under three sensing and visualization conditions: (A) Baseline (Unaided Vision), (B) AR + LiDAR Detection, and (C) AR + Multi-Headset Sharing. For each of the two scenarios, 30 trials were performed under each of the three conditions, resulting in a total of 180 outdoor trials. The sample size of 30 trials per condition was selected to ensure stable mean estimates based on the central limit theorem. Time to collision (TTC) was used as the primary metric for evaluating early hazard perception.
Across trials, start positions and approach trajectories were held constant. Inherent differences in walking pace, reaction time, and approach speed introduced trial to trial variability. Participants were instructed to walk or drive at a comfortable pace without specific speed targets. In the pedestrian–e-scooter scenario, pedestrian walking speed averaged 1.2 m/s (SD = 0.15) and e-scooter speed averaged 3.2 m/s (SD = 0.4). In the pedestrian–vehicle scenario, pedestrian speed averaged 1.1 m/s (SD = 0.12) and vehicle speed averaged 6.5 m/s (SD = 0.8). Conditions were counterbalanced across sessions to mitigate learning and fatigue effects.
@@ IV-C Experimental Scenarios
Scenario 1: Pedestrian–E-Scooter Interaction. A pedestrian and an e-scooter rider approached a shared intersection from orthogonal directions. The e-scooter emerged from a fully occluded 90 corner at 10–15 km/h, while the pedestrian walked toward the expected intersection point (EIP) at normal pace as shown in Figure 2.
Scenario 2: Pedestrian–Vehicle Interaction. This scenario involved a pedestrian crossing near a side road while a vehicle approached from a different direction. The vehicle emerged from an approximately 80 curved corner at 20–30 km/h, gradually entering the pedestrian’s path of motion. Figure 3 presents the layout of this real-world interaction.
@@ IV-D Experimental Conditions
Condition A: Unaided Eye (Baseline). Neither the pedestrian nor the counterpart (e-scooter rider or vehicle driver) wore an AR device. Situational awareness depended entirely on natural human vision. Ground truth positions and velocities for TTC computation were obtained from VLP-16 LiDAR trajectories.
Condition B: AR + LiDAR Detection. The pedestrian wore an MQPro AR headset, while the counterpart did not. The counterpart’s position and velocity were estimated using the VLP-16 LiDAR, and the pedestrian received real-time AR cues (3D bounding boxes and directional arrows) corresponding to the detected hazards.
Condition C: AR + Multi-Headset Sharing. Both the pedestrian and the counterpart wore MQPro headsets. SLAM-based multi-headset alignment provided a unified world coordinate frame, enabling direct AR to AR exchange of headset poses without relying on LiDAR. Although occasional SLAM resets occurred during fast vehicle motion, the overall tracking remained sufficiently stable for TTC analysis.
IV-E Time To Collision Computation
@@
$\mathrm{TTC}=\frac{d_{\mathrm{EIP}}}{v},$ (13)
@@
TABLE I: Time to Collision (TTC) across different sensing and visualization configurations for two real-world near collision scenarios.
@@ IV-F Summary of Results
Table I presents the average TTC values computed across 30 repeated trials for each configuration, quantitatively demonstrating that ARCAS substantially improves early hazard perception across both scenarios. Several key findings emerge. Across all conditions, TTC values for the pedestrian increased substantially under AR guidance. In the pedestrian–e-scooter scenario, TTC increased from 2.75 s in the Unaided Eye baseline to 5.51 s with AR + LiDAR and 5.43 s with multi-headset sharing, indicating an approximately twofold improvement in the pedestrian’s available reaction time. The e-scooter rider also exhibited substantial gains, with TTC increasing from 0.69 s to approximately 2.86 s under AR conditions, more than a fourfold improvement. In the pedestrian–vehicle scenario, AR + LiDAR Detection produced the highest pedestrian TTC (6.42 s), while multi-headset Sharing delivered a similar improvement (6.29 s) despite occasional SLAM resets. Overall, these results confirm that ARCAS, whether powered by external LiDAR sensing or AR to AR SLAM sharing, provides significantly earlier hazard awareness and larger safety margins for VRUs in occluded and fast approaching traffic conditions.
@@ V Conclusion
This paper introduced ARCAS, a real-time Augmented Reality Collision Avoidance System designed to enhance VRUs’ situational awareness in dynamic traffic environments. By integrating 360° 3D LiDAR sensing, SLAM-based headset tracking, and automatic 3D calibration, the system delivers world locked visual cues, such as 3D bounding boxes and directional arrows, directly in the user’s passthrough view. ARCAS further supports multi-headset alignment, enabling multiple VRUs to share a unified world frame for coordinated awareness. From a broader perspective, ARCAS represents a step toward an intersection-level digital twin; however, a full digital twin implementation would require additional components such as a persistent world model, and scenario replay capabilities, which remain directions for future development. The system was evaluated through two real-world near collision scenarios involving pedestrian interactions with an e-scooter rider and a vehicle. Across 180 trials, AR visualization consistently increased TTC for both the pedestrian and the counterpart. In the pedestrian–e-scooter scenario, the pedestrian’s TTC nearly doubled and the rider’s TTC increased more than fourfold under AR guidance. In the pedestrian–vehicle scenario, LiDAR-based detection yielded the greatest improvements, while multi-headset sharing achieved similar gains despite occasional SLAM resets. These results demonstrate that AR-based visual cues, whether from external sensing or AR to AR pose sharing, provide substantially earlier hazard awareness and larger reaction windows in occluded or fast approaching conditions.
Overall, ARCAS establishes a practical foundation for AR-based safety systems that deliver real-time, personalized hazard awareness for VRUs. Future work will explore intent prediction, adaptive cue generation, and context aware filtering, with expanded evaluation using additional surrogate safety measures such as post encroachment time and deceleration rate to avoid collision. Future studies will also address human factors including user acceptance, cognitive workload, and trust in wearable AR for traffic safety. Moreover, the designed current experiments were conducted in a controlled campus environment, and future work should evaluate ARCAS in more complex and dense urban traffic settings to advance intelligent, human-centered safety solutions for next generation mobility systems.